### 18 Dec 2023 v2

# RecJPQ: Training Large-Catalogue Sequential Recommenders

Abstract.
Sequential Recommendation is a popular recommendation task that uses the order of user-item interaction to model evolving users’ interests and sequential patterns in their behaviour. Current state-of-the-art Transformer-based models for sequential recommendation, such as BERT4Rec and SASRec, generate sequence embeddings and compute scores for catalogue items, but the increasing catalogue size makes training these models costly. The Joint Product Quantisation (JPQ) method, originally proposed for passage retrieval, markedly reduces the size of the retrieval index with minimal effect on model effectiveness, by replacing passage embeddings with a limited number of shared sub-embeddings. This paper introduces RecJPQ, a novel adaptation of JPQ for sequential recommendations, which takes the place of item embeddings tensor and replaces item embeddings with a concatenation of a limited number of shared sub-embeddings and, therefore, limits the number of learnable model parameters. The main idea of RecJPQ is to split items into sub-item entities before training the main recommendation model, which is inspired by splitting words into tokens and training tokenisers in language models. We apply RecJPQ to SASRec, BERT4Rec, and GRU4rec models on three large-scale sequential datasets. Our results showed that RecJPQ could notably reduce the model size (e.g., 48$\times$ reduction for the Gowalla dataset with no effectiveness degradation). RecJPQ can also improve model performance through a regularisation effect (e.g. +0.96% NDCG@10 improvement on the Booking.com dataset). Overall, RecJPQ allows the training of state-of-the-art transformer recommenders in industrial applications, where datasets with millions of items are common.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 17th ACM International Conference on Web Search and Data Mining; March 4–8, 2024; Merida, Mexicobooktitle: Proceedings of the 17th ACM International Conference on Web Search and Data Mining (WSDM ’24), March 4–8, 2024, Merida, Mexicoprice: 15.00doi: 10.1145/3616855.3635821isbn: 979-8-4007-0371-3/24/03
1. Introduction
111
Sequential recommender systems are a class of recommendation models that use the sequence of user-item interactions to predict the next item. Most of the state-of-the-art models for sequential recommendation are based on deep neural networks, for example, recurrent neural networks (Hidasi et al., 2016; Hidasi and Karatzoglou, 2018), convolutional neural networks (Tang and Wang, 2018; Yuan et al., 2019), and most recently, transformers (Kang and McAuley, 2018; Sun et al., 2019; Petrov and Macdonald, 2022a, b). All these models use learnable item embeddings as an essential component in their model architectures. Figure 1 illustrates item embeddings in a typical neural sequential recommendation model. As the figure shows, item embeddings usually have two roles in the model architecture: (i) to convert the sequence of input item ids to a sequence of item representation vectors and (ii) to convert the sequence embedding produced by the model into the distribution of predicted item scores. In both cases, a recommender system that works with an item set $I$ requires an embedding tensor with $|I|\cdot d$ parameters, where $d$ is the size of each embedding.
222
When a recommender has many items in the catalogue, various challenges arise in training the neural recommendation model. Firstly, the item embedding tensor may contain more model parameters than the rest of the model. For example, there are more than 800 million videos on YouTube (you, 2023). If a recommender model uses 128-dimensional embeddings, the whole item embeddings tensor will have more than 100 billion parameters, which is comparable with the number of parameters of the largest available machine learning models (Brown et al., 2020), even without accounting for the parameters of the model’s intermediate layers. This is a problem specific to recommender systems: in the related area of dense passage retrieval (Karpukhin et al., 2020; Khattab and Zaharia, 2020), passage embeddings are obtained by encoding passage text using a pre-trained language model; however, item side information, such as text, is not necessarily available in a typical recommender systems scenario; therefore, item embeddings should be directly learned from the interactions. Secondly, a large number of such trainable parameters also makes the model prone to overfitting. A third challenge caused by the large catalogue is the size of the output scores tensor (rightmost tensor in Figure 1): for example, in BERT4Rec, it contains a score for each item for every position, for every sequence in the training batch, so training BERT4Rec with more than 1 million items in the catalogue may be prohibitively expensive (Petrov and Macdonald, 2022a). This problem is typically solved using negative sampling, whereby instead of computing the full output tensor, the model computes scores for a small proportion of negative items not interacted by the user (e.g. SASRec (Kang and McAuley, 2018) uses one negative per positive). However, negative sampling comes with its own challenges (for example, it usually requires informative negative mining (Rendle, 2022)). Nevertheless, negative sampling is an orthogonal research direction, and in this paper, we use SASRec in cases where negative sampling is necessary. We refer the reader to our recent publication (Petrov and Macdonald, 2023), where we analysed the negative sampling problem in details. To summarise the challenges, a large embedding tensor increases the model size, slows model training down, and can necessitate further modelling tricks such as negative sampling, which bring their own challenges.
There are some existing methods (Xia et al., 2023; Wang et al., 2022; Kang et al., 2020) for item embedding compression (we discuss these methods in Section 2). However, most of these methods compress the embedding tensor after the model is fully trained (including training the full embedding tensor). However, as argued above, training may be prohibitively expensive in large-catalogue recommender systems. Hence, this paper addresses the problem of a large item embedding tensor in sequential recommendation models at the training stage.
To mitigate this problem, we propose a novel RecJPQ technique inspired by the success of a recent Joint Product Quantisation (JPQ) work (Zhan et al., 2021a) for passage retrieval. JPQ itself is based on Product Quantisation (PQ) (Jégou et al., 2011), a popular method of compressing vectors by splitting them into sub-embeddings and encoding them using a discrete item [^3] For consistency, we explain prior work using item ids instead of passage ids. codebook (the codebook maps from item ids to the associated sub-item ids; see Section 3.1 for the details). The main innovation of JPQ compared to the standard PQ method is that it learns the sub-item embeddings as part of the overall model training process. In contrast, PQ requires training the model first and only then compressing the embeddings (frequently, this second step uses external tools, such as FAISS (Johnson et al., 2021)). This means that JPQ does not need to keep the embedding matrix in memory during model training. We argue that this innovation is valuable for recommender systems. Indeed, as mentioned above, real-life recommender systems can have hundreds of millions of items in their catalogues and keeping full embeddings tensor in memory may be prohibitively expensive. This is particularly important for deep-learning-based sequential recommender systems because these models require keeping the whole model in GPU (or TPU) memory during training. GPU memory is costly even when compared to regular computer RAM.
Figure 1. Item embeddings in a typical sequential recommender system. These item embeddings are used in two ways: (i) to obtain sequence representation and (ii) to generate item scores. The embedding tensor requires $|I|\cdot d$ trainable parameters, where $|I|$ is the items catalogue size, and $d$ is the size of an embedding. When catalogue size $|I|$ is large, item embeddings comprise most of the model’s parameters.
444
Unfortunately, it is hard to adapt JPQ to the recommendation scenario, as it is specific to textual information retrieval. In particular, JPQ assumes the existence of a pre-trained (passage retrieval) model and index, which it uses to assign items to sub-item ids (see more details in Section 3.2). These pre-trained models rarely exist in item recommendations. Hence, in RecJPQ, we experiment with performing the initial assignment of sub-ids using three different strategies: (i) discrete truncated SVD (sub-ids obtained by discretising the item representations obtained by an SVD decomposition of the user-item matrix), (ii) discrete BPR (sub-ids obtained by discretising the item embeddings obtained from BPR) and (iii) random assignments. We describe these assignment strategies in detail in Section 4.
RecJPQ is a model component that replaces traditional item embeddings in sequential recommender systems. In general, it can be applied to any recommender system based on item embeddings, but in this paper, we focus specifically on sequential models, as in these models, item embeddings comprise the biggest part of the model (e.g. sequential models usually do not have user embeddings). In contrast with existing methods, RecJPQ does not require training full uncompressed embedding and does not modify the original model loss function. Our experimentation on three datasets (see Section 5) demonstrates that RecJPQ can be successfully applied to different models, including SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019) and GRU (Hidasi et al., 2016; Petrov and Macdonald, 2022a), achieving a large factor of embeddings compression (e.g. 47.94$\times$ compression of SASRec on Gowalla) without any effectiveness degradation. Moreover, on 2 out of our 3 experimental datasets, applying RecJPQ increases model performance (e.g. +0.96% NDCG@10 on Booking.com dataset, significant improvement); we attribute these improvements to model regularisation.
In short, the contributions of this paper are as follows: (1) we propose RecJPQ, a novel technique for reducing the size of sequential recommendation models during training based on Joint Product Quantisation; (2) we propose three strategies for assigning sub-item ids to items, two of which (discrete truncated SVD and discrete BPR) assign similar codes to similar items, and one assigns codes randomly; (2) we perform an extensive experimental evaluation of RecJPQ on three datasets and show that RecJPQ allows reducing the models’ size without hindering the model performance. The rest of the paper is organised as follows: Section 2 introduces related work on embeddings compression and identifies limitations of existing methods; Section 3 covers Product Quantisation (PQ) and Joint Product Quantisation (JPQ) – the methods, which serve as the basis for our work; Section 4 introduces RecJPQ and covers sub-id assignment strategies for RecJPQ; in Section 5 we experimentally evaluate RecJPQ; Section 6 contains final remarks.
555
2. Related work
Table 1. Existing embedding compression methods. The desired method characteristics are highlighted in bold.
Model Agnostic Method Backbone Sequential backbone Trains full embeddings Compression unit
No EODRec (Xia et al., 2023) SASRec (Kang and McAuley, 2018) Yes Yes Item
LightRec (Lian et al., 2020) DSSM (Huang et al., 2013) No Yes Item
MDQE (Wang et al., 2022) SASRec (Kang and McAuley, 2018) Yes Yes Item
Yes PreHash (Shi et al., 2020a) BiasedMF (Koren et al., 2009); NeuMF (He et al., 2017) No No User
Quotient Remainder (Shi et al., 2020b) DCN (Wang et al., 2017); DLRM (Naumov et al., 2019) No No Features
MGQE (Kang et al., 2020) SASRec (Kang and McAuley, 2018); NeuMF (He et al., 2017); GCF (He et al., 2017); Yes Yes Item
Yes RecJPQ (ours) SASRec (Kang and McAuley, 2018); BERT4Rec (Sun et al., 2019); GRU (Hidasi et al., 2016) Yes No Item
This section covers existing work on compressing and discretising embeddings in recommender systems, identifies the limitations in existing work and positions our contributions in the context of existing methods. Table 1 summarises existing methods and positions RecJPQ, our proposed compression technique. The table highlights with boldface the desirable characteristics necessary for training a large-scale [^6] For simplicity, we say that a catalogue is ”large-scale” if it has more than 1 million items, as it becomes challenging to train recommender systems on that scale (Petrov and Macdonald, 2022a). sequential model, specifically: the method can be applied to work with different backbone sequential models, and does not require training full embeddings (as we work with the assumption that the full embeddings tensor does not fit into GPU memory); and we want the model to focus on item embeddings rather than embeddings of other entities, such as users or features. As illustrated in the table, the methods for compressing the models can be broadly divided into two groups: model dependent and model agnostic.
In the model-dependent methods (Lian et al., 2020; Xia et al., 2023), the embedding compression mechanism is integrated as a component into the recommendation model itself. Hence, the training architecture of these methods has to be aware of the compression, and the loss function includes components responsible for the embedding compression. For example, LightRec (Lian et al., 2020) uses the Deep Semantic Similarity Model (DSSM) (Huang et al., 2013) as the backbone model and uses an additional knowledge distillation component in the loss, which allows for learning compressed representations of the embeddings. However, DSSM is not a sequential model, and it is unclear whether or not the method can be adapted to the sequential recommendation task. Similarly, EODRec (Xia et al., 2023), uses SASRec (Kang and McAuley, 2018) as its backbone, one of the most popular sequential models based on the Transformer architecture (Vaswani et al., 2017). The loss function of EODRec also consists of four components, some of which are responsible for recommendation, and others are responsible for embedding compression. While EODRec’s backbone (SASRec) is an efficient model, in many cases, other models such as BERT4Rec exhibit better results (Sun et al., 2019; Petrov and Macdonald, 2022b). In general, while some model-dependent methods may reduce the size of a trained model, these methods have several limitations, which make them unsuitable for training sequential recommendation models with large catalogues. In particular:
L1 Model-dependent methods are, by their nature, tied to a specific model, making them inflexible when adapting to a specific task. For example, LightRec (Lian et al., 2020) uses a non-sequential DSSM model as a backbone. The core component of LightRec (Recurrent Composite Encoding) is tightly integrated into the DSSM architecture, and it is unclear whether or not it can be used outside of DSSM.
777
L2 Model-dependent methods usually require training (uncompressed) item embeddings and then use knowledge distillation or teacher-student techniques to obtain compressed representations of the embeddings. This approach substantially reduces the final model size, thereby helping inference on smaller devices, but requires a large amount of GPU memory while training, thereby limiting the overall number of items in the catalogue. For this reason, the main positioning for EODRec model (Xia et al., 2023) is the on-device recommendation: while the final model produced by this method is small, it requires storing full item embeddings while training. Post-training quantisation methods (Yao et al., 2022), which recently became popular to reduce the size of large language models via quantising their weights into lower-precision numbers (e.g. float16, or int8) also have this limitation – they need to have access to the full model before quantising. Similarly, Mixed Precision Training (Micikevicius et al., 2018) builds a smaller precision model, but it requires keeping full precision weights in memory. Placing the embeddings tensor into Approximate Nearest Neighbours (Johnson et al., 2021) or Hierarchical Navigable Small Worlds (Malkov and Yashunin, 2020) indexes also requires access to the full embedding tensor at model training time, and therefore also exhibits this limitation.
L3 Model-dependent methods require multi-component loss functions, some of which are responsible for the recommendation task and others for the model compression. This is a form of multi-objective optimisation, which is a challenging problem (Sener and Koltun, 2018; Deb and Deb, 2014), as finding the balance between the loss components for different objectives usually requires extensive hyperparameters search.
On the other hand, the existing model-agnostic methods (Shi et al., 2020a, b) do not depend on the specific model architecture, and likewise do not add extra components to the loss functions. Typically, these methods implement a mechanism that takes the place of the embeddings tensor in the backbone model, and hence can be used with many models. However, on inspection of the relevant work, we elicit additional limitations of these methods:
L4 Many methods are not designed for compressing item embeddings. For example, PreHash (Shi et al., 2020a) is a method specific for compressing user embeddings (i.e. it uses the user’s history to construct user embeddings). The method uses an attention network over the history of user interactions. Adapting this network structure for items embeddings is a hard task: a user may only interact with a few items; in contrast, a popular item may be interact with by millions of users. The attention mechanism depends quadratically on the sequence length, and therefore, applying it to users who interacted with a popular item would be prohibitively expensive.
L5 Finally, many methods lack structure in compressed item representations. This leads to situations where unrelated items have similar representations and, conversely, when similar items have dissimilar representations. Both these cases may limit the models’ generalisation ability and hinder the models’ performance. One example of such unstructured methods is hashing-based Quotient Remainder (Shi et al., 2020b), which compresses embeddings of categorical features (e.g., genres). Quotient Remainder assigns feature codes based on the quotient and the remainder of the division of the feature id by some number. When applied to item ids (items can also be seen as categorical features), the quotient and the remainder are unrelated to the item characteristics. Hence, similar items are unlikely to have similar codes. Nevertheless, Quotient Remainder is one of the few methods that can be used to train a model on a large-scale dataset, and therefore, we use this method as a baseline in our experiments.
888
Overall, among the related work, we argue that existing methods exhibit a number of Limitations (L1-L5). In summary, the model-dependent methods require training full embeddings first, limiting the maximum number of items that can be considered in the catalogue of the recommender system. On the other hand, methods which do not require training full embeddings first, such as Quotient Remainder, rely on heuristics and may result in unrelated items having similar representations. On the surface, the nearest related work to ours is VQ-Rec (Hou et al., 2023) as it also applies JPQ-style technique to sequential recommendation; however, similarly to JPQ, it relies on the availability of textual features and pre-trained language models. In contrast, our work’s main novelty is adapting JPQ to the scenario when (e.g., textual) side information is not available. In the next section, we describe Product Quantisation – a vector compression method and JPQ (Zhan et al., 2021a), a method of embedding compression for information retrieval, which directly learns embeddings in compressed form, reducing GPU memory requirements. Then, in Section 4, we propose RecJPQ – an adaptation of JPQ to the sequential recommendation task, which successfully addresses Limitations L1-L5.
Table 2. Analysis of PQ’s impact on memory requirements for storing item embeddings tensor for selected recommendation datasets, based on 512-dimension float32 vector embeddings. The table compares base memory usage with different code lengths, shown as percentages relative to the base.
Dataset Num Items Size of Item Embedding Tensor
Base Code length=2 512 centroids 1.00 MB Code length=8 2,048 centroids 4.00 MB Code length=32 8,192 centroids 16.00 MB
MovieLens-1M 3,416 6.67 MB 14.988% 59.953% 239.813%
Booking.com 34,742 67.86 MB 1.474% 5.895% 23.580%
Gowalla 1,280,969 2.44 GB 0.040% 0.160% 0.640%
3. Product Quantisation and JPQ
We now describe Product Quantisation (PQ) and Joint Product Quantisation (JPQ), two methods that serve as backbones for our method. Section 3.1 covers PQ, a classic embedding compression technique. Section 3.2 describes JPQ – a recently proposed information retrieval method that learns compressed embeddings directly instead of compressing them after the model training.
3.1. Product Quantisation
Product quantisation (Gray, 1984; Jégou et al., 2011) is a well-cited method of compressing vectors used by many libraries, such as FAISS (Johnson et al., 2021) & nanopq (Yusuke Matsui, 2023). Its main idea is to split a collection of $d$-dimensional vectors, $V$, into $m$ collections $V_{i};i=1..m$ of smaller vectors of $\frac{d}{m}$ dimensions each. The original vectors can be recovered back via concatenation: $V=concat(V_{1},V_{2},...V_{m})$. Product quantisation then clusters each $V_{i}$ into $b$ clusters (e.g. using the k-means algorithm (MacQueen, 1967)) and replaces each vector $v_{ij}$ with the centroid of the assigned cluster $c_{ij}\approx v_{ij}$. With this replacement, the original vectors collection can be approximated as a concatenation of the sub-vector matrices $C_{1}...C_{m}$, which are constructed from $V_{1}...V_{m}$ by replacing each vector $v_{ij}$ by the closest cluster centroid $c_{ij}$: $V\approx concat(C_{1},C_{2},...C_{m})$. In each sub-vector matrix $C_{i}$, there are, at most, $b$ different rows, as each row corresponds to one of the centroids of the clusters, so instead of storing full matrix $C_{i}$, we can store these unique rows in the separate tensor $\hat{C}_{i}$, whose elements $\hat{c}_{ij},j\in\{1..b\}$ correspond to the unique sub-vectors. To compress the approximate embedding matrix, we need only store the ids of the sub-vectors (sub-ids) for each item. Overall, there are $m$ sub-vector sets $C_{i}$, so each vector can be encoded using $m$ integer codes, and each code can have $b$ different values; therefore, overall, this scheme can encode up to $b^{m}$ different vectors. The vector of sub-item ids $g_{i}=\{g_{i1},g_{i2}...,g_{im}\}$ associated with the vector $v_{i}\in V$ is known as the code of the vector $v_{i}$ (Jégou et al., 2011). The number $m$ of sub-item ids associated with each item is called the length of the code. The table $G$ of codes associated with each vector from $V$ is also known as a codebook (Jégou et al., 2011).
Figure 2 illustrates the vector reconstruction process applied to item embeddings. For each sub-item id $g_{ij}$ in the codes vector $g_{i}$ of an item $i$, we extract a sub-item embedding associated with this sub-item id and then concatenate the sub-item embeddings to obtain reconstructed item embeddings.
Figure 2. Reconstruction of item embeddings using Product Quantisation: Codebook length $m=3$, item embedding length $d=12$, number of sub-ids per split $b=256$.
999
The number of splits $m$ is usually considerably smaller than the original vector dimensions $d$: $m\ll d$ to achieve compression. The number of sub-embeddings per split, $b$, is typically a power $k$ of 256, so that the codes can be stored as $k$-byte integers. In this paper, for simplicity and following JPQ (Zhan et al., 2021a), we fix $k=1$, so each sub-item id can be represented with a single byte; therefore, we only store $m$ bytes for each item. Even with fixed $k$ (and therefore $b$), we can adjust model capacity by controlling the number of sub-item ids associated with each item, $m$, and the dimension of the sub-item embeddings, $\frac{d}{m}$, by adjusting $d$. Figure 2 illustrates PQ applied to item embeddings for $m=3$, $d=12$, $b=256$. Further, to illustrate the achieved compression, if embeddings are stored as 256-dimensional float32 vectors, a full-item embedding requires 1KB of memory. After compression using product quantisation with $m=8$ splits, we only need to store 8 bytes per item (0.78% of the original memory requirement). While some memory is also required to store the sub-item embeddings themselves, this is negligible for large datasets compared to the original vector requirements (see also Table 2 for an analysis of sub-item embeddings memory requirements).
Product Quantisation is a well-established vectors compression technique, which has been shown to be successful in approximate nearest neighbour methods (Jégou et al., 2011; Johnson et al., 2021; Ge et al., 2014; Li et al., 2020), information retrieval (Khattab and Zaharia, 2020; Santhanam et al., 2022; Izacard et al., 2020) and recommender systems (Balen and Levy, 2019; Kang et al., 2020; Wang et al., 2022). However, Product Quantisation requires the full embeddings tensor to be trained before the embeddings are compressed. Indeed, the quantisation operation is not differentiable, and therefore, the model can not be trained end-to-end. Therefore, it requires first training full (non-quantised) model and only then applying compression. Some recommendation models (e.g. (Kang et al., 2020)) use differentiable variations of Product Quantisation to allow end-to-end training, but these methods still require training full embeddings alongside the quantised versions.
Overall, Product Quantisation addresses Limitations LABEL:limitation:model_tied (it is model agnostic), LABEL:limitation:not-items (it is applicable for training item embeddings compression), and LABEL:limitation:not-similar (when sub-item ids are assigned using clusterisation, similar items will have similar codes). However, it does not address Limitations LABEL:limitation:full_embedding (it requires training full embeddings first), and LABEL:limitation:complex_loss (PQ uses a separate loss function for embeddings reconstruction, which is not aligned with ranking loss). In the next section, we discuss Joint Product Quantisation, a method that can be adapted to addressing these remaining limitations.
3.2. Joint Product Quantisation
Zhan et al. recently proposed Joint Product Quantisation (JPQ) (Zhan et al., 2021a), a Product Quantisation-based method developed for dense information retrieval. The main difference with the classic product quantisation is that JPQ generates item codes before training the model. As code assignment is the only non-differentiable operation in Product Quantisation, therefore, when the codes are assigned before training the model, the model can be trained end-to-end without training full item embeddings – JPQ essentially replaces the embeddings tensor in the model, where item embeddings are constructed via sub-item embeddings concatenation, as illustrated in Figure 2. Assuming that the codebook in the figure is a constant, all other parameters can be learned using standard gradient descent on model’s loss. In particular, this means that JPQ does not require special loss function components to learn sub-item embeddings. 101010 Compared to the original Product Quantisation method, JPQ addresses Limitations LABEL:limitation:full_embedding (it does not require training full item embeddings) and LABEL:limitation:complex_loss (it does not require a specific loss function). However, in contrast to plain PQ, JPQ does not provide a mechanism to assign similar embeddings to similar items and, therefore, does not address Limitation LABEL:limitation:not-similar. The sub-item id assignments method proposed in the original JPQ paper is specific for text retrieval (as it relies on the existence of a pre-built index for a text document collection generated using the STAR model (Zhan et al., 2021b)). In the next section, we introduce RecJPQ, an adaptation of JPQ to the sequential recommendation scenario, which does not rely on text-specific datasets and models. Our adaptation of JPQ to sequential recommendation requires careful design of novel sub-item id assignment strategies. Indeed, to the best of our knowledge, this is the first adaptation of the JPQ method to sequential recommendation.
4. RecJPQ
RecJPQ is a Joint Product Quantisation-based method for training recommendation models with a large catalogue of items. As we discuss in Section 3.2, the method used by JPQ for initial sub-item id assignments relies on the existence of a pre-built index of documents and, therefore, can not be directly used for recommendation scenarios. Hence, the main difference between RecJPQ and the original JPQ is sub-item id assignment strategies. In general, RecJPQ can be described as performing the following steps:
Step 1. Build the item-code mapping matrix (codebook) using one of the sub-item id assignment strategies (see Section 4.1).
Step 2. Initialise the sub-item embeddings randomly.
Step 3. Replace the item embedding tensor with the concatenation of sub-item embeddings associated with each item).
Step 4. Train the model end-to-end using the model’s original training task and loss function (this process also trains the sub-item embeddings, so they do not need to be trained separately).
The way RecJPQ builds the codebook before training the main model is also similar to how language models (such as BERT (Devlin et al., 2019)) train a tokenisation algorithm before training the main model. Language models also learn embeddings of sub-word tokens instead of learning embeddings of full words to reduce the model’s size; similarly, RecJPQ learns sub-item embeddings instead of learning embeddings of full items. Below, we describe sub-item id assignment strategies used by RecJPQ.
111111
4.1. Sub-Item Id Assignment Strategies
Random Sub-Item Id Assignments. In the most simple scenario, we can assign items to sub-item ids randomly. We compose the item code out of $m$ random integers in this case. RecJPQ with random sub-item id assignments strategy does not address Limitation LABEL:limitation:not-similar (similar items do not have similar codes). Indeed, with random sub-item id assignments, RecJPQ becomes similar to other ”random” embeddings compression methods, such as the hashing trick (Weinberger et al., 2010) or the Quotient Reminder method (Shi et al., 2020b). The main problem with these methods is that unrelated methods are forced to share parts of their representation, which limits the generalisability of the models. However, as we show in Section 5, sometimes random assignments may be beneficial, as the random assignments strategy acts as a form of regularisation. Nevertheless, as the random sub-item id assignments strategy does not address Limitation LABEL:limitation:not-similar (similar items should have similar representations), we introduce further sub-item id assignment strategies that can address this limitation in the next sections.
Discrete Truncated SVD. As discussed in Section 3.2, the only limitation not addressed by JPQ is Limitation LABEL:limitation:not-similar (similar items should have similar codes). Random sub-item id assignments, as discussed above, do not address this limitation either. Hence, in this section, we design a sub-item id assignments method that addresses this remaining limitation and assigns similar codes to similar items. Some approaches have used side information, such as textual data for item representations (Rajput, Shashank et al., 2023); however, we address the more generic classic sequential recommendation scenario and, therefore, must infer item similarities from the user’s sequences. To do this, we employ the SVD algorithm, which has been shown to achieve good results in learning item representations (Zhou et al., 2012) for recommender systems.
We first compute a matrix of sequence-item interactions $M$, where rows correspond to sequences (users), and columns correspond to items. This matrix’s elements $m_{ij}$ are either 1 if $i^{th}$ contains interactions with item $j$ and 0 otherwise. We then compute the truncated SVD decomposition of matrix $M$ with $m$ latent components: $M\approx U\times\Sigma\times V^{T}$, where $U$ is the matrix of user embeddings, $V$ is the matrix of item embeddings, and $\Sigma$ is the diagonal matrix of largest singular values.
Our initial experiments showed that some items have equal embeddings after performing truncated SVD decomposition. This occurs when two items have been interacted with by exactly the same set of users. To ensure that all items have different embeddings, we normalise $V$ using the min-max normalisation range and add a small amount of Gaussian noise: $\hat{v}_{ab}=\frac{v_{ab}-\min_{k}{v_{ak}}}{\max_{k}{v_{ak}}-\min_{k}{v_{ak}}}+\mathcal{N}(0,10^{-5});\forall v_{ab}\in{V}$ The variance of the noise ($10^{-5}$) is negligible compared to the range of possible values of normalised embeddings ($[0..1]$ after min-max normalisation). Therefore it has a very small influence on the position of the items in the embeddings space. However, if two items have exactly the same embeddings after decomposition (e.g. this can happen if two items appear in exactly the same set of sequences), the noise allows us to distinguish these two embeddings. Lastly, the assignment of sub-item id involves discretising each dimension of the normalised item embeddings into $b$ quantiles so that each quantile contains an approximately equal number of items. We use these bins as sub-item id assignments for the items. Note that although this method requires computing an $m$-dimensional item embeddings tensor (and there can be hundreds of millions of items), it does not require computing them as part of a deep learning model training on a GPU. Indeed, truncated SVD is a well-studied problem. There are effective algorithms for performing it that do not require modern GPUs (Halko et al., 2011). Moreover, as the method only requires computing $m$-dimensionsional embeddings, the table will be many times smaller than full $d$-dimensional embeddings, so the method requires $\frac{d}{m}$ times less memory to store embeddings. Finally, performing truncated SVD is possible in a distributed manner (e.g., using Apache Spark (spa, 2023)), which allows truncated SVD to be applied on large datasets. 121212 In summary, discrete truncated SVD allows assigning similar codes to similar items; it does not require a GPU for intermediate computations and can be easily performed for very large datasets.
Discrete BPR. Truncated SVD is not the only Matrix Factorisation method that can be used for initial sub-item id assignments. In particular, we also use the classic BPR approach (Rendle et al., 2009) to obtain coarse item embeddings. The method also learns user embeddings (or, in our case, sequence embeddings) $U$ and item embeddings $V$. The estimate of the relevance of an item $i$ for user $j$ is defined as the dot product of user and item embeddings: $r=u_{j}\cdot v_{i}$. In contrast with truncated SVD, BPR does not directly approximate the user-item interaction matrix. Instead, BPR optimises a pairwise loss function that aims to ensure that positive items are scored higher than negative items: $\mathcal{L}_{BPR}=-\log(\sigma(u_{i}\cdot v_{j^{+}}-u_{i}\cdot v_{j^{-}}))$, where $v_{j^{+}}$ is the embedding of a positive item for the user $u$, $v_{j^{-}}$ is the embedding of a randomly sampled negative item, and $\sigma$ is the sigmoid function.
BPR is one of the most cited methods in recommender systems; therefore, we use BPR as an alternative strategy for coarse item embedding learning. The rest of the discrete BPR strategy is the same as in the truncated SVD: we also normalise the learned embeddings using a min-max normalisation and add a small amount of Gaussian noise to ensure different embeddings for different items are obtained. Similar to truncated SVD, BPR does not require learning on a GPU, and there exist distributed implementations (Luke de Oliveira, 2018) that allow for learning item embeddings on very large datasets, so overall, it can be used as an alternative to truncated SVD for sub-item id assignments in RecJPQ.
This concludes the description of the sub-item id assignment strategies for RecJPQ. We now discuss why RecJPQ may act as a regularisation mechanism and improve the performance on the datasets with many long-tail items.
4.2. RecJPQ as a Regularisation Mechanism
The interactions with items in recommender systems typically exhibit a long tail distribution (Park and Tuzhilin, 2008), meaning that a few popular items receive the most interactions. In contrast, most items comprise the ”long tail” with few interactions. As the training data for these long-tail items is limited, recommendation models can suffer from overfitting on such items (Zhang et al., 2021), causing overall performance degradation.
Goodfellow et al. (Goodfellow et al., 2016, Ch. 7) argued that one of the most powerful regularisation techniques is parameter sharing – a technique where certain parameters of the model are forced to be equal. RecJPQ is a special case of parameter sharing: we force different items to share parts of their embeddings. This prevents the model from learning embeddings that are too specific to only a few training sequences, as each part of the embedding appears in many other sequences via the sharing mechanism. In our experiments (Section 5), we indeed observe that RecJPQ may act as a model regulariser and improve the model’s performance; this is especially apparent in the Gowalla dataset, where the proportion of long-tail items is the largest.
RecJPQ Summary. To conclude Section 4, we summarise that RecJPQ is a model component that takes the place of the item embeddings tensor in sequential recommender systems. RecJPQ is based on the JPQ method, which is a variation of the PQ method. RecJPQ addresses all of the limitations described in Section 2: it is model-agnostic (Limitation LABEL:limitation:model_tied); does not require training full embeddings (Limitation LABEL:limitation:full_embedding); does not modify the backbone model’s loss function (Limitation LABEL:limitation:complex_loss); it is suitable for item embeddings compression (Limitation LABEL:limitation:not-items); it can assign similar codes to similar items with the help of discrete truncated SVD or discrete BPR (Limitation LABEL:limitation:not-similar). Futhermore, we argue that RecJPQ may act as a model regulariser, which is an additional advantage when there are many long-tail items in the catalogue. This concludes the description of RecJPQ. In the next section, we experimentally evaluate RecJPQ and analyse its effects on required memory and on the model performance.
131313
5. Experiments
Table 3. Characteristics of the datasets. ”Long tail %” is the percentage of items with less than 5 interactions.
Dataset Users Iems Interactions Avg. length Long tail %
MovieLens-1M 6,040 3,416 999,611 165.49 0.0%
Booking.com 140,746 34,742 917,729 6.52 61.8%
Gowalla 86,168 1,271,638 6,397,903 74.24 75.8%
Our experiments address the following research questions:
  1. RQ1
    What is the effect of the sub-item id assignment strategy?
  2. RQ2
    How do code length $m$ and embedding size impact effectiveness?
  3. RQ3
    What is the effect of RecJPQ on size/effectiveness tradeoff?
5.1. Experimental Setup
Backbone Models. In our experiments, we use two state-of-the-art Transformer-based recommendation models: BERT4Rec (Sun et al., 2019) – a model that uses a transformer encoder based on BERT (Devlin et al., 2019); and SASRec (Kang and McAuley, 2018) – a model which utilises decoder part of the Transformer (similar to GPT (Radford et al., 2019)). For both models, we use the versions [^14] The code for the paper is available at https://github.com/asash/RecJPQ from our recent reproducibility paper (Petrov and Macdonald, 2022b), which provides efficient & effective implementations (using the Huggingface Transformers library (Wolf et al., 2020)). Additionally, to demonstrate that RecJPQ can be applied to other architectures, we use a GRU (Cho and van Merrienboer, 2014)-based model from (Petrov and Macdonald, 2022a) available in the same repository. This model uses the GRU4Rec (Hidasi and Karatzoglou, 2018) architecture, but a slightly different configuration, e.g. it uses LambdaRank (Burges, 2010) as a loss function, which is shown to be effective (Petrov and Macdonald, 2022a).
Datasets. We experiment with three datasets: (i) MovieLens-1M (denoted ML-1M) (Harper and Konstan, 2015) – this is a movie rating dataset that is one of the most popular benchmarks for recommender systems; (ii) Booking.com (Goldenberg and Levin, 2021) – a multi-destination trips dataset, and (iii) Gowalla (Cho et al., 2011) – a check-in dataset. Following common practice (Sun et al., 2019; Petrov and Macdonald, 2022b), we remove users with less than five interactions. Table 3 lists the salient characteristics of the datasets after preprocessing. As can be seen from the table, the number of items in these datasets varies from relatively small (3416 in MovieLens-1M) to large (1,271,638 in Gowalla) – this allows testing RecJPQ in different settings (RecJPQ is designed for large datasets, and we expect it to compress the model by a larger factor on Gowalla).
Table 4. Impact of RecJPQ with different sub-item id assignment strategies on model size and effectiveness for the MovieLens-1M and Booking.com datasets. Relative Size corresponds to model checkpoint size as the percentage of the base model. ${}^{=}$, ${}^{+}$, and ${}^{-}$ denote significance testing results compared to the base, respectively: indistinguishable ($pvalue>0.05$, Bonferroni multi-test correction), better or worse. Bold denotes the best NDCG@10 in each column.
[…]
ML-1M
Base 0.157 100.0% 0.072 100.0% 0.131 100.0%
Hashing (Quotient-Remainder) ${}^{-}$ 92.4% ${}^{-}$ 61.6% ${}^{-}$ 124.9%
RecJPQ-BPR ${}^{=}$ 93.2% ${}^{=}$ 62.5% ${}^{=}$ 128.0%
RecJPQ-Random ${}^{=}$ 93.2% ${}^{=}$ 62.5% ${}^{=}$ 127.6%
RecJPQ-SVD ${}^{=}$ 93.2% ${}^{=}$ 62.5% ${}^{=}$ 127.9%
Booking
Base 0.376 100.0% 0.209 100.0% 0.137 100.0%
Hashing (Quotient-Remainder) ${}^{-}$ 62.8% ${}^{-}$ 27.6% ${}^{-}$ 9.2%
RecJPQ-BPR ${}^{=}$ 63.3% ${}^{+}$ 27.5% ${}^{+}$ 8.7%
RecJPQ-Random ${}^{-}$ 62.3% ${}^{+}$ 27.5% ${}^{+}$ 8.9%
RecJPQ-SVD ${}^{+}$ 63.3% ${}^{+}$ 27.6% ${}^{+}$ 8.8%
The datasets are also diverse regarding the number of “long-tail items” (defined as items with less than five interactions). While the MovieLens-1M dataset does not have long-tail items, the Booking.com dataset has 60.8% long-tail items, and Gowalla has 75.8% long-tail items. As discussed in Section 4.2, RecJPQ acts as a model regulariser in long-tail distributions, and we expect to see the highest regularisation effect on the Gowalla dataset.
Evaluation Protocol. Overall, our evaluation protocol follows the protocol from the recent replicability paper (Petrov and Macdonald, 2022b). We use a leave-one-out data splitting strategy: we hold out the last sequence in each sequence in the test set. Additionally, for 1024 randomly selected users, we hold out the second last action into a validation set, which we use for the early stopping mechanism. We set the maximum sequence length at 200. If the sequence contains more than 200 interactions, we use 200 latest interactions. If the sequence contains less than 200 interactions, we left-pad it to ensure its length is exactly 200. To ensure that the models are fully converged, following (Petrov and Macdonald, 2022a), we employ an early stopping mechanism on the NDCG@10 metric: we stop training if the metric is not improved for 200 epochs.
Metrics. Our main focus is the trade-off between model size and model effectiveness. For measuring effectiveness, following prior research (Kang and McAuley, 2018; Sun et al., 2019; Petrov and Macdonald, 2022b), we use NDCG@10, and as the model size metric, we use the file size of the model checkpoint. Following best practices (Cañamares and Castells, 2020; Krichene and Rendle, 2022; Dallmann et al., 2021), we measure NDCG without using negative sampling.
151515
Baselines. We deploy an adaptation of Quotient Remainder (Shi et al., 2020b) as a baseline compression approach, applied to each base model – this parameter-free hashing-based approach encodes each item using two hashes: the quotient and the remainder of the division of item id by $\left\lceil\sqrt{|I|}\right\rceil$ where $|I|$ is the catalogue size. Quotient Remainder guarantees that each item has a unique code.
We do not apply post-training embedding quantisation (e.g. float16), nor use other methods from Table 1 as baselines, as they are not suitable for our task: EODRec, LightRec, MDQE and MGQE require training full embeddings (we assume that training full embeddings is not an option for a large catalogue), and PreHash is specific for compressing user embeddings, so is not suitable for item embeddings. However, reducing the model size by decreasing the embedding dimensionality can also be seen as a simple baseline. We analyse models using different embedding sizes in RQ3.
To analyse the effect of the sub-item id assignment strategy on model performance/model size tradeoff, we compare the original (base) versions of BERT4Rec, SASRec and GRU with RecJPQ versions trained with Random, discrete truncated SVD and discrete BPR sub-item id assignment strategies. We do not train GRU and BERT4Rec on Gowalla, as these models do not use negative sampling. Training models on this dataset without negative sampling is not feasible due to the large GPU memory requirement for storing output scores (Petrov and Macdonald, 2022a), while applying negative sampling is a substantial change to the models’ training process that is outside of the scope of this paper. In all cases, we use 512-dimensional embeddings and the code of length $m=8$ (we experiment with other embedding sizes and lengths of the code in the next section). One exception is for the SASRec base model on Gowalla; in this case, we use 128-dimensional item embeddings (item embeddings larger than 128 dimensions consume all available GPU memory when embedding compression techniques are not deployed).
5.2. Results
RQ1. Effect of sub-item id assignment strategy. Table 4 shows the experimental results on the smaller ML-1M and Booking datasets, while Table 5 reports results for the Gowalla dataset. The tables compare NDCG@10 and model size of compressed variations of backbone models with the base (uncompressed) model. Significant differences compared to the corresponding base model (BERT4Rec, GRU or SASRec) are indicated. In general, the tables show that RecJPQ substantially reduces the model checkpoint size in most cases. For example, the RecJPQ versions of the GRU models on the Booking dataset are approximately 27% of the original in size. On the Gowalla dataset, compressed models are approximately 3% of the original. Moreover, model size does not depend on the sub-item id assignment strategy. Indeed, sub-item id assignments only influence the values of the model parameters but not the number of parameters. Moreover, Quotient Remainder models have approximately the same compression level as RecJPQ models. We speculate that after compression, the model checkpoint size is dominated by other model parameters (e.g., attention matrices). In our configuration, the sub-item embeddings tensor only requires a few megabytes of memory (see Table 2). In contrast, the full model checkpoint of a compressed model is typically tens of megabytes (e.g., 92.8MB for SASRec using RecJPQ-BPR trained on Gowalla).
161616
RecJPQ only increased the model size for SASRec on MovieLens-1M, due to the dataset’s small item count: on this dataset, the overhead of storing sub-item embeddings and the codebook is larger compared to the benefit of compressing the embeddings table. Using RecJPQ with smaller embeddings might reduce the model size without affecting effectiveness on this dataset (see also RQ3).
On the other hand, we observe from Table 4 and Table 5 that the choice of the best assignment strategy depends on both the model and the dataset. For example, on MovieLens-1M, the choice of the strategy is not important, and in all cases, RecJPQ versions of the models are statistically indistinguishable from all corresponding base models. On the larger Booking dataset, the choice of the best strategy is model-dependent. For BERT4Rec, the best results are achieved with BPR (NDCG@10 0.375, statistically indistinguishable from the base) and SVD (NDCG@10 0.379, +0.97%, significant). At the same time, the Random strategy significantly underperforms the base configuration (NDCG@10 0.316, -15.98%) – this shows that in some cases, assigning similar codes to similar items is indeed important. However, in 2 cases, Random performs statistically significantly better than SVD and BPR. For example, Random assignments perform best on Gowalla with the SASRec base (a significant improvement of +57% over the base). SVD assignments also moderately improve the result in this case (+10%, significant). At the same time, BPR decreases the quality by a large margin on Gowalla dataset (-70%) [^17] The percentages for the Gowalla dataset seem large because this dataset is difficult: it has the largest number of items and the largest proportion of long-tail items.. We explain the success of the Random strategy on the Gowalla dataset as giving a larger regularisation effect (random assignments make the learning task harder, so the model has fewer chances to overfit). Overall, from Tables 4 and 5, we conclude that there is no ”one size fits all” choice of the sub-item id assignment strategy, and it depends on both dataset and model salient characteristics; the exact best combination of model/strategy may depend on regularisation requirements (as we observe in Gowalla), the prevalence of strong sequential patterns (as in Booking, see (Petrov and Macdonald, 2022a) for details) and so on. This suggests that the sub-item id assignment strategy could be treated as a hyperparameter and tuned for each model/dataset combination. However, by default, we recommend using RecJPQ with the SVD strategy – in all cases, this achieves significantly better (on the Booking and Gowalla datasets) or statistically indistinguishable (on the MovieLens-1M dataset) results compared to the base model. We also note that RecJPQ with the SVD strategy is always better than the Quotient Remainder baseline (Quotient Reminder is always significantly worse than the base model, whereas RecJPQ is better or indistinguishable). The question of whether or not it is possible to select the best model/strategy combination without doing an exhaustive hyperparameter search is an interesting research direction, which we leave for future work.
Table 5. Impact of RecJPQ with different sub-item id assignment strategies on SASRec model size and effectiveness on the large-scale Gowalla dataset. Notations follow Table 4.
Strategy NDCG @10 Relative Size
Base 0.110 100.0%
Hashing (Quotient-Remainder) ${}^{-}$ 2.8%
RecJPQ-BPR ${}^{-}$ 2.8%
RecJPQ-Random ${}^{+}$ 2.9%
RecJPQ-SVD ${}^{+}$ 2.9%
(a) MovieLens-1M
(b) Gowalla
Figure 3. RecJPQ performance for SASRec while varying embedding size $d$ and the number of sub-item ids per item $m$.
181818
It is also worth mentioning that we do not observe a degradation of the training efficiency when training the RecJPQ versions of the models. Indeed, while there are some fluctuations in the training time the model requires to converge, the magnitude of the required training time remains the same: for example, training of the base version of BERT4Rec requires 18.8 hours on Booking.com, and the RecJPQ-SVD version of BERT4Rec requires 16.1 hours. The training time of the SVD model (used for initial sub-item id assignments) in the same case is negligible compared to the training of the main model (it takes approximately a minute). Inference time is also unaffected (e.g. on Gowalla, full evaluation across the 86k users requires 10 minutes for both ”base” and RecJPQ versions of the models).
In summary, in answer to RQ1, we conclude that RecJPQ achieves large model compression levels. The compression is particularly impressive on datasets with large catalogues (like Gowalla). The compression does not depend on the sub-item id assignment strategy. However, the sub-item id assignment strategy greatly affects the model performance. The effect is model- and dataset-dependent, so the strategy should be treated as a hyperparameter. However, the SVD is a safe choice, as it always provides results that are comparable (i.e statistically indistinguishable) or better than the base model.
RQ2. Effects of code length $m$ and the embedding size on model performance. To answer RQ2, we perform a grid search over embedding size and code length on the MovieLens-1M and Gowalla datasets. We use SASRec as the backbone (the only model that can be easily trained on Gowalla) and apply the SVD sub-item id assignment strategy. We select the embedding size $d$ from {$2^{0},2^{1},...,2^{9}$} and code length $m$ from {$2^{0},2^{1},...,2^{8}$}. Note that $m\leq d$, as RecJPQ splits each embedding of size $d$ into $m$ sub-embeddings.
191919
Figure 3 illustrates the results of the grid search. The figure shows the NDCG@10 of the SASRec-RecJPQ model for each combination of code length (x-axis) and embedding size (y-axis), in the form of a heatmap for both datasets. As we can see from the figure, a larger embedding size generally positively affects the model performance. This result echoes similar findings of a recent reproducibility paper (Rendle et al., 2022); however, interestingly, in the RecJPQ case, increasing embedding dimensionality does not change the amount of information we store per each item, as the length of the code defines it rather than the embedding size. Instead, it increases model capacity, increasing the amount of information that can be stored in each sub-item embedding, allowing to account for more item characteristics. For example, on Gowalla, the largest embedding we can train using base SASRec is 128 dimensions, while with RecJPQ, we can train the model even with 512-dimensional embeddings.
On the other hand, larger code lengths are not always helpful. As we discussed in Section 4.2, RecJPQ forces the model to share parts of embeddings with other items and, therefore, acts as a regularisation mechanism. A shorter code length forces items to share more information, and therefore, it causes a stronger regularisation effect. As we can see, on the less sparse MovieLens-1M – where all items have more than five interactions – regularisation is not an issue, and longer codes are beneficial. For example, the best result is achieved with 512-dimensional embeddings and a code of length 128 (NDCG@10 0.14). In contrast, for Gowalla, where most items are long-tail items with less than five interactions (hence the embeddings of these items should be regularised), the best NDCG is achieved with the code of length 8 (NDCG@10 0.12). The fact that the model can perform better with shorter codes confirms that RecJPQ can behave as a regularisation technique.
202020
In short, in answer to RQ2, we conclude that larger embeddings are generally beneficial for model performance. However, the sparser Gowalla dataset benefits from shorter code lengths, due to the regularisation effect of parameter sharing brought by RecJPQ.
(a) MovieLens-1M
(b) Gowalla
Figure 4. NCDG/Size tradeoff for SASRec and SASRec-RecJPQ.
RQ3. Size-Performance tradeoff. To address our last research question, we analyse the trade-off between model checkpoint size and NDCG@10 achieved by the model when trained with different embedding sizes. We select the embedding size from {1, 2, 4, 8, 16, 32, 64, 128, 256, 512} and train the original versions of SASRec and SASRec-RecJPQ with the SVD strategy on the MovieLens-1M and Gowalla datasets. For RecJPQ, we select code length $m$ that is optimal for the dataset/embedding size pair (according to the grid search from Figure 3). For the original SASRec on Gowalla, we only train up to the embedding size of 128 due to GPU memory limit.
Figure 4 illustrates the tradeoff between model checkpoint size and NDCG@10 for both SASRec and SASRec-RecJPQ on the two datasets. Each point on the figure corresponds to one embedding size. As can be seen from the table, a larger model size (corresponding to larger embeddings) leads to better performance for both SASRec and SASRec-RecJPQ (this echoes findings in the previous research question). However, SASRec-RecJPQ’s performance grows much faster with increasing model size than observed for vanilla SASRec. For example, the largest vanilla SASRec model achieves roughly the same performance as the 4.9$\times$ smaller SASRec-RecJPQ version of the model (71MB vs. 15 MB). This effect is even more prominent in Gowalla, where the number of items is larger: the largest SASRec model achieves roughly the same performance as the 47.94$\times$ smaller SASRec-RecJPQ model (3.2GB vs. 69MB).
Overall, in answer to RQ3, we conclude that while larger models benefit model performance, RecJPQ improves this tradeoff by a large margin (i.e. to achieve the same performance, RecJPQ requires much fewer parameters than the original model). This effect is more markedly pronounced for the larger Gowalla dataset.
6. Conclusions
In this paper, we discussed the challenge of training sequential recommender systems with large datasets, primarily due to the large item embedding tensor. Existing embedding compression methods have limitations, leading to our proposed method, RecJPQ. Our evaluation of RecJPQ on three datasets resulted in significant model size reduction, e.g., 47.94$\times$ compression of the SASRec model on the Gowalla dataset. Additionally, RecJPQ serves as a model regulariser, improving the model’s quality, with SASRec-RecJPQ using SVD strategy outperforming the original SASRec model (+35% NDCG@10 on Booking, +10% on Gowalla). There are a number of future work directions, which can be based on this paper, e.g., an analysis of the RecJPQ effect on training/inference time and analysing the theoretical grounds for the observed regularisation effects.